Please help me solve this equation: 1 2/3 - 2 1/2
Found 2 solutions by checkley77, leighm:
Answer by checkley77(12844)

1 2/3 - 2 1/2
5/3-5/2
(5*2-5*3)/6
(10-15)/6
-5/6 answer.

Answer by leighm(1)

To do this problem, it is easiest to start off my making the numbers have the same denominator. So 1 2/3 = (3/3)+(2/3) = 5/3
and 2 1/2 = (4/2)=(1/2) = 5/2. But we need a common denominator, so multiply 3*2 to find the common denominator 6. The multiply 5/3 by 2 (since 3*2=6), and 5/2 times 3 (since 2*3=6). Add the remaining fractions.
So 10/6 + 15/6 = 25/6